OpenStudy (anonymous):

2 questions will medal please help!!!

OpenStudy (anonymous):

What are the questions?

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Jackson purchased a pack of game cards that was on sale for 22% off. The sales tax in his county is 6%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards.

OpenStudy (anonymous):

y times .22 = x y-x = final cost

OpenStudy (anonymous):

22*6

OpenStudy (anonymous):

oh sorry i didnt see the sales tax part haha

OpenStudy (anonymous):

you can do what i did above, but then factor in sales tax by that number by .06 and adding that to the final price

OpenStudy (anonymous):

First I would put 22% and 6% into a decimal. You can do this by dividing the numbers by 100.

OpenStudy (anonymous):

n Martins class of 20 students, 4 students report that social studies is their favorite class and 75% of the students report that PE is their favorite class. Show your work to explain how to determine whether more students prefer social studies or PE.

OpenStudy (anonymous):

When we use sale rates, we multiply the number (in this case the variable y) by the percent that's taken off (22% or .22 as a decimal) When we get that number we subtract them together then multiply that number by sales tax (6% or .06 as a decimal) when we get this number we then subtract it into how much the item cost.

OpenStudy (anonymous):

4/20 = 20%

OpenStudy (anonymous):

15/20 = 75%

OpenStudy (anonymous):

I would say the answer would be : y (.22 - .06) or y (.22.^.06)